OpenStudy (anonymous):

factor out the GCF from the expression -6x^6+30x^5-18x^4. can someone help me with this problem?

OpenStudy (anonymous):

You could use euclid's algorithm to find the GCD of 6, 18, and 30 which is 6. The smallest power is 4, so you would factor out 6x^4

OpenStudy (cwrw238):

the gcf is the greatest number or variable which divides exactly in the number/variable so gcf of -6,30 and 18 = 6 gcf of x^6, x^4 and x^5 = x^4 reqd. gcf = 6x^4

OpenStudy (anonymous):

is this the final answer

OpenStudy (cwrw238):

do you just want the gcf or the complete factors?

OpenStudy (anonymous):

it says factor out the gcf from the expression

OpenStudy (cwrw238):

hhm not sure to get the other factors you divide 6x^4 into each term eg -6x^6 / 6x^4 = -x^(6-4) = -x^2 so starting you off 6x^4( -x^2 + - ) can you fill in the blanks?

OpenStudy (anonymous):

hello

OpenStudy (cwrw238):

lol hi!

OpenStudy (anonymous):

no help me

OpenStudy (anonymous):

I think the answer your question is looking for is 6x^4 Do you know euclid's algorithm?